One angle of a triangle is a right angle. Another angle of the triangle is three times the smallest angle. What are the measures of all the angles?
Answer by John10(297)

The total degrees of the triangle is 180 degrees.
Since there is a right triangle, it is 90 degrees
So the other two angles are 90 degrees
Let x(smallest angle) and y be those two angles: x + y = 90 (1)
Another angle is three times the smallest number: y = 3x (2)
Substitute (2) into (1)
x + 3x = 90
4x = 90
x = 22.5
Substitute x = 22.5 into (2) to find y:
y = 3(22.5)= 67.5
The angles are 22.5, 67.5, and 90.
Hope it will help you!John10